Suppose a study is being planned to estimate the relative frequency of companies that are labeled as risky. What sample size is needed so that the standard error will be no larger than 0.016 ?
Answer #1

here as we know that maximum standard error is at p=0.5

therefore required sample size =p(1-p)/(standard error)2 =0.5*(1-0.5)/(0.016)^2 =976.56 ~ 977